OpenStudy (anonymous):

A triangle has two angles measuring 78º and 43º . What is the measure of the third angle?

OpenStudy (anonymous):

would it be 121

OpenStudy (anonymous):

In a triangle the sum of the angles is 180°, so 180=78+43+x (It can't possily be 121)

OpenStudy (anonymous):

A triangle has three angles and all the three of the them add upto 180 degrees (in any damn triangle)

OpenStudy (anonymous):

So, 180=78+43+third angle 78+43=121 So, 180=121+third angle 180-121=third angle 59= third angle

OpenStudy (anonymous):

@╰☆╮Openstudier╰☆╮ I know it seems redundant in most cases, but try to hint the answer rather than give it away in the future :)

OpenStudy (anonymous):

Fine @Quantaliinuxite